OpenStudy (anonymous):

completely simplify (x-2)/(8x^2- 32)

OpenStudy (anonymous):

\[\frac{x-2}{8x^2-32}\]

OpenStudy (anonymous):

(x-2)/(8(x^2-4))

OpenStudy (anonymous):

(x-2)/(8(x+2)(x-2)) \[\frac{ (x-2) }{ (8(x+2)(x-2)) }\]

OpenStudy (anonymous):

first factor out an 8 out of the bottom \[\frac{x-2}{8(x^2-4)}\] next use \[(a-b)^2 =(a-b)(a+b)\] when a and b are perfect squares

OpenStudy (anonymous):

cancel the common now

OpenStudy (anonymous):

\[a^2-b^2=(a+b)(a-b)\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

1/8(x+2) is this right?

OpenStudy (anonymous):

= (x-2)/8(x^2-4) = (x-2)/8(x+2)(x-2) = 1/8(x+2)

OpenStudy (anonymous):

yes thats absolutely corret:)

OpenStudy (anonymous):

Thank you